OBSERVATION OF INTERSTELLAR H3^{+}

\underline{BENJAMIN J. McCALL^{a}}, Department of Chemistry, The University of Chicago, Chicago, IL 60637; KENNETH H. HINKLE, National Optical Astronomy Observatories, Tucson, AZ 85726; THOMAS R. GEBALLE, Joint Astronomy Centre, Hilo, HI 96720; TAKESHI OKA, Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL 60637.

The long awaited initial detection^{b} of H3^{+} in the interstellar medium provides unprecedented opportunites to directly characterize the physical and chemical properties of dense molecular clouds near young stellar objects. Recent observations in the 3.7 μm region using the CGS4 infrared spectrometer at the United Kingdom Infrared Telescope and the highresolution Phoenix spectometer at the Kitt Peak National Observatory support the initial detection of H3^{+} and extend the use of H3^{+} as a chemical probe of the interstellar medium to other molecular clouds. The interstellar material towards many winter objects — GL961E, MonR2 IRS 2 and IRS 3, GL490, GL989, NGC2024 IRS 2, LkHα 101, W3 IRS 5, and Orion BN — have been searched in the spectral region of H3^{+}, providing both positive (detection) and negative (upper limit) results. In addition, the data from the initial detection of H3^{+} in the summer objects GL2136 and W33A have been reprocessed to improve the resolution of the spectral lines.

aSupported by the Fannie and John Hertz Foundation

bT.R. Geballe and T. Oka, Nature \underline{384}, 334 (1996)